# On the Girth of Graph Lifts

### v1

@@ 3 Moore and Erdös Sachs Bounds for Graph Lifts
The proof consists of two parts. The first part is the existence of some graph $G$ covering $H$ with girth at least $g$, no matter how large $G$ is. The second part is showing that as long as $\mbox{diam}(G)$ is larger than the bound above, $G$ can be trimmed into a smaller graph $G^{\prime\prime}$ covering $H$ without introducing cycles shorter than $g$. We prove the first part in Lemma 6 and the second part here. We may assume that $G$ is connected, because otherwise we take $G^{\prime\prime}$ as one of its connected components.
@@
Let $G$ be be some graph of girth $g\geq g_{0}$ covering $H$ with the cover map $\pi:G\rightarrow H$ and assume that there are vertices $u^{\prime},v^{\prime}\in V(G)$ with $\mbox{dist}(u^{\prime},v^{\prime})>R$. Let $u=\pi(u^{\prime})$, $v=\pi(v^{\prime})$, where possibly $u,v$ are the same vertex. Without loss of generality we identify $V(G)$ with $V(H)\times[n]$ and $E(G)$ with $E(H)\times[n]$ so that the $n$ copies of $T$ in $\pi^{-1}(T)$ respect layers. Namely, that for any edge $e\in E(T)$, both endpoint of each edge in $\pi^{-1}(e)$ are in the same layer. Furthermore, since $\mbox{dist}(v^{\prime},u^{\prime})>R>\mbox{diam}(T)$ it is clear than $v^{\prime}$ and $u^{\prime}$ are not in the same layer and we may assume that $v^{\prime}=(v,n)$ and that $u^{\prime}=(u,n-1)$.
Let $L_{i}$ denote the vertices at layer $i$ and let $L_{[j]}=L_{1}\cup\cdots\cup L_{j}$. We define $G^{\prime}$ as the induced graph of $G$ on $L_{[n-2]}$ and the map $\pi_{G^{\prime}}$ as the restriction of the cover map $\pi:G\rightarrow H$ to $G^{\prime}$. Let $\cal{R}$ be the set of edges $e\in E(G)$ with tail in $L_{n-1}\cup L_{n}$ and head in $L_{[n-2]}$. Note that $\pi_{G^{\prime}}$ is not a cover map because of the edges in $\cal{R}$, which we call red edges, leave deficient vertices in $G^{\prime}$, i.e. vertices $w\in V(G^{\prime})$ with $\deg(w)<deg(\pi(w))$. Since edges covering $E(T)$ respect layers, it follows that red edges are necessarily non-tree edges covering $E(H)\setminus E(T)$. Moreover, for any non-tree edge $e\in E(G)$, at most one of its endpoints is in $L_{n-1}\cup L_{n}$. Indeed, if both endpoint are in the same layer $L_{i}$, then we have a short cycle using $e$ and the shortest path in the tree $\pi^{-1}(T)\cap L_{i}$ whose length is at most $1+\mbox{diam}(T)<g_{0}$. Otherwise, if $t(e)=(x,n)$ and $h(e)=(y,n-1)$ then we have a short path from $v^{\prime}=(v,n)$ to $u^{\prime}=(u,n-1)$ of length at most $d(v^{\prime},t(e))+d(t(e),h(e))+d(h(e),u^{\prime})\leq 2\mbox{diam}(T)+1\leq R$. As both cases lead to contradiction, it follows there are exactly two red edges above any non-tree edge $e\in E(H)\setminus T$, which we denote by $((t(e),n),(h(e),l(e)))$ and $((t(e),n-1),(h(e),m(e)))$.
@@
Figure 1 depicts the red edges above some non-tree edge in $e\in E(H)\setminus E(T)$ and its inverse. In case (a) $e=(x,y)$ is not its own inverse, and therefore it has 4 distinct red edges, two above $e$ and two above $e^{-1}$. In case (b) $e=(z,z)$ is equal to its inverse and it has only two red edges.
@@
In order to prove that $\mbox{girth}(G^{\prime\prime})$ is at least $g$, assume to the contrary that $G^{\prime\prime}$ has a length $k<g$ cycle $C$ with the edges $e_{1},e_{2},\ldots,e_{k}$. Since $\mbox{girth}(G^{\prime})\geq g$, the cycle $C$ has at least one new (blue) edge. We say that the new edge $e$ is positive if $t(e)$ is connected in $G$ by a red edge to $L_{n-1}$ and $h(e)$ is connected to $L_{n}$. Otherwise, $t(e)$ is connected to $L_{n}$ and $h(e)$ to $L_{n-1}$ and we say that $e$ is negative. If all new edges in $C$ have the same sign, then there is a path in $G$ of length at most $k-1+2=k+1$ connecting $L_{n}$ with $L_{n-1}$ and therefore $d_{G}(v^{\prime},u^{\prime})\leq 2\mbox{diam}(T)+k+1\leq 2\mbox{diam}(T)+g$, contradicting our assumption that $\mbox{dist}_{G}(v^{\prime},u^{\prime})>R$.
Otherwise, there are at two new edges $e,f$ in $C$ with a different sign. Therefore there is a path $P$ of length at most $\lfloor(k-2)/2\rfloor$, consisting of old edges from $h(e)$ to $t(f)$. Since both $h(e)$ and $t(f)$ are connected by red edges in $G$ to the same layer, $L_{n}$ or $L_{n-1}$, it follows that $G$ has a cycle of length at most $\lfloor(k-2)/2\rfloor+2+\mbox{diam}(T)\leq k/2+1+\mbox{diam}(T)<g/2+1+\mbox{diam}(T)$, which is less equal than $g$ for $g\geq g_{0}$, yielding a contradiction. ∎
Lemma 6.
[…]
@@
The proof is by induction on $g$, where the base $g=1$ trivially holds for $G=H$. So, assume that $G$ covers $H$ with $\mbox{girth}(G)=g-1$ and with the minimal possible number of length $g-1$ cycles, $\phi(G)>0$.
Let $\cal G$ be a random 2-lift of $G$. Namely, a random graph with the vertex set $V(G)\times\{0,1\}$, where for each edge $e\in\overline{E}(G)$ we pick the permutation $\mbox{perm}_{\pi}(e)\in S_{2}$ independently at random such that $\mbox{perm}_{\pi}(e)=\mbox{perm}_{\pi}(e^{-1})$. As all lifts of $G$ have girth at least $g-1$, our task is to find a 2-lift with a smaller number of length $g-1$ cycles. Since every length $g-1$ cycle $C$ in $G$ is turned into either one cycle of length $2(g-1)\geq g$ or two cycles of length $g-1$, depending on the product of $\mbox{perm}_{\pi}(e)$ for $e\in C$, it follows that $E[\phi(\cal G)]=\phi(G)$. However, since the lift with $\mbox{perm}_{\pi}(e)=\mbox{id}$ for all $e$ has $2\phi(G)$ length $g-1$ cycles, it follows that there exists some 2-lift of $G^{\prime}$ with $\phi(G^{\prime})<\phi(G)$, yielding a contradiction. ∎
@@ 4 Three Generalizations for Non-regular Graphs
Respectively, $n(H,g)=n({\cal L}(H),g)$ is the smallest girth $g$ lift of $H$, $n(\tilde{H},g)=n({\cal U}(H),g)$ is the smallest girth $g$ graph with the same universal cover as $H$ and $n({\cal D}(H),g)$ is the smallest girth $g$ graph with the same degree distribution as $H$. Since ${\cal L}(H)\subseteq{\cal U}(H)\subseteq{\cal D}(H)$ it follows that:
@@
    $\frac{1}{|\overline{E}|}\sum_{e\in\overline{E}(H)}||(B_{H})^{r}\cdot\delta_{e}||_{1}\geq\Lambda^{r}$ for all $r\geq 0$, where $\delta_{e}$ is the characteristic vector of $e$.
@@
This result is related to the $\rho$ based lower bound of Theorem 1 by the following Lemma:
Lemma 9.
For any finite undirected and connected graph $H$ we have
$\rho(H)\geq\Lambda(H).$
Proof.
Given some graph $H$, we know that $\rho(\tilde{H})=\lim_{r\rightarrow\infty}|B_{r}(v)|^{1/r}$ for any $v\in V(\tilde{H})$. Therefore, for every $\epsilon>0$ there exists $r_{0}$ so that $\max_{v}|B_{r}(v)|\leq(\rho(H)+\epsilon)^{r}$ for every $r\geq r_{0}$. Let $G$ be a girth $g=2r_{0}+1$ covering graph for $H$. Then by Claim 8, $(\rho(H)+\epsilon)^{r}\geq\max_{v}|B_{r}(v)|\geq\Lambda^{r}$, yielding the required result. ∎
It follows from Claim 8, Lemma 9, inequalities (6), (5) and Theorem 1 that:
@@
It should be observed that because all three graphs are non-regular, we have strict inequality $\Lambda(H)>\overline{d}(H)-1$, in agreement with Claim 8. Regarding the comparison of $\Lambda(H)$ and $\rho(H)$, the situation is more interesting:
@@
We have no high level argument to justify the strict inequality $\rho(H_{23})>\Lambda(H_{23})$, which basically says that the growth rate of $\widetilde{H_{23}}$, show in Figure 3 (b), is strictly higher that what is obtained by the Markov chain argument of AHL2002 [1]. See our conjecture at the end, that $\rho(H)>\Lambda(H)$ whenever the non-backtracking random walk is irreducible and aperiodic.
@@ 5 Concrete Constructions and Bounds for the $H_{23}$ Graph
In this section we explore non-asymptotic upper and lower bounds on the value of $n(H,g)$, for the graph $H=H_{23}$ depicted in Figure 2 (b). This we do in the following subsections: The Moore lower bound, The Erdös Sachs Upper Bound, Construction Algorithms, and Results.
@@ 5.1 The Moore Lower bound
The Moore lower bound for lifts is stated in Lemma 3, where the ball size around some tree vertex can be obtained by summing (4) on $r$. Therefore, for any $g=2r+1$ we have:
$n(H,g)\geq n_{0}(H,g)=\max_{v\in V(\tilde{H})}|B_{r}(v)|=1+\max_{v\in V(H)}\sum_{i=0}^{r-1}||\left(B_{H}\right)^{i}\delta(v)||_{1},$ (9)
where $\delta(v)$ is the indicator function of $t^{-1}(v)$. One can derive a similar formula for even girth.
5.2 The Erdös Sachs Upper Bound
We compute the Erdös Sachs upper bound for $n(H,g)$ using Corollary 5. As the spanning tree $T$ has diameter one, it follows that for all $g\geq g_{0}=4$:
$n(H,g)\leq\min_{v\in V(\tilde{H})}|B_{g+2}(v)|=1+\min_{v\in V(H)}\sum_{i=0}^{g+1}||\left(B_{H}\right)^{i}\delta(v)||_{1}.$
@@ 5.4 Result Summary
6 Conclusion
In this work we generalize the well studied problem of finding $n(k,g)$, which is the smallest size of a $k$-regular girth $g$ graph. Given some fixed base graph $H$, we define: (i) $n({\cal L}(H),g)=n(H,g)$ as the smallest size of a girth $g$ graph covering $H$, (ii) $n({\cal U}(H),g)=n(\tilde{H},g)$ as the smallest size of a girth $g$ graph with the same universal cover as $H$, and (iii) $n({\cal D}(H),g)$ as the smallest size of a girth $g$ graph with the same degree distribution as $H$. It should be noted that only the second and third definitions offer a true generalization of $n(k,g)$, by taking any $k$-regular graph as $H$.
We prove that the first two definitions are equivalent up to a multiplicative constant, and we show that the well known Moore lower bound and Erdös Sachs upper bound hold in the new setting as well, with the exponent base being $\rho(H)$ rather than $k-1$. For the 3rd definition we have a Moore bound result by Alon, Hoory and Linial [1]. However, there is no Erdös Sachs result and the Moore bound might be far from achievable as we conjecture is the case for $H_{23}$.
@@ 6 Conclusion
    Conjecture: For any graph $H$, with minimal degree at least two, we have $\rho(H)=\Lambda(H)$ only if $H$ is: (i) Regular, or (ii) bi-regular, or (iii) a regular or bi-regular graph where each edge is replaced by a length $k+1$ path.
    Stated differently, we conjecture that Lemma 9 is a strict inequality if the non-backtracking random walk is irreducible.
@@
    Conjecture: $n({\cal D}(H_{23}),g)\geq{\rho(H_{23})}^{(1+o(1))g/2}$
    The conjecture states that the exponent base lower bound on the size of girth $g$ graphs with 50% degree 2 and 50% degree 3 vertices can be improved from $\Lambda(H_{23})$ to $\rho(H_{23})$. One way to prove it is to show that such minimal size graphs cannot have adjacent degree 2 vertices.
@@
    Problem: Prove that any of the greedy algorithms listed in Section 5 satisfies the asymptotic Erdös-Sachs upper bound of Theorem 1.
@@
    Problem: Prove some version of the following: There is no non-trivial Moore graph with girth $g>g_{0}$, where $g_{0}$ is some constant and a Moore graphs is defined as a graph satisfying (9) as an equality, possibly up to a divisibility constraint as we have seen in the $H_{23}$ case.
  6. 6.
    Problem: Give a number theoretic construction of girth $g$ graphs covering some non-regular graph $H$ of size ${\rho(H)}^{(1+o(1))\cdot 3g/4}$.

### v2

@@ 3 Moore and Erdös-Sachs Bounds for Graph Lifts
Given some base graph $H$ and spanning tree $T$ satisfying the requirements, denote $g_{0}=2\mbox{diam}(T)+2$ and $D_{0}=g+2\mbox{diam}(T)$. The proof consists of two parts. The first part is to prove the existence of some graph $G$, possibly very large, covering $H$ with girth at least $g$. The second part is showing that as long as $\mbox{diam}(G)>D_{0}$, one can be trim $G$ into a smaller graph covering $H$ without introducing a cycle shorter than $g$. Repeating this procedure until $\mbox{diam}(G)\leq D_{0}$ yields the required result. We prove the existence of an initial graph $G$ separately in Lemma 3.4, and concentrate here of the trimming procedure. Throughout, we assume that $G$ is connected, as otherwise it can be replaced by one of its connected components.
@@
Let $G$ be be some graph of girth $g\geq g_{0}$ covering $H$ with the cover map $\pi:G\rightarrow H$ with vertices $u^{\prime},v^{\prime}\in V(G)$ such that $\mbox{dist}(u^{\prime},v^{\prime})>D_{0}$. Let $u=\pi(u^{\prime})$, $v=\pi(v^{\prime})$, where possibly $u,v$ are the same vertex. Without loss of generality we identify $V(G)$ with $V(H)\times[n]$ and $E(G)$ with $E(H)\times[n]$ so that the $n$ copies of $T$ in $\pi^{-1}(T)$ respect layers. Namely, that for any edge $e\in E(T)$, both endpoint of each edge in $\pi^{-1}(e)$ are in the same layer. Furthermore, since $\mbox{dist}(v^{\prime},u^{\prime})>D_{0}>\mbox{diam}(T)$ it is clear than $v^{\prime}$ and $u^{\prime}$ are not in the same layer, and we may assume that $v^{\prime}=(v,n)$ and $u^{\prime}=(u,n-1)$.
[…]
@@
Figure 1 depicts the red edges above some non-tree edge in $e\in E(H)\setminus E(T)$ and its inverse. In case (a) $e=(x,y)$ is not its own inverse, and therefore has 4 distinct red edges, two above $e$ and two above $e^{-1}$. In case (b) $e=(z,z)$ is equal to its inverse and it has only two red edges.
@@
In order to prove that $\mbox{girth}(G^{\prime\prime})$ is at least $g$, assume to the contrary that $G^{\prime\prime}$ has a length $k<g$ cycle $C$ with the edges $e_{1},e_{2},\ldots,e_{k}$. Since $\mbox{girth}(G^{\prime})\geq g$, the cycle $C$ has at least one new (blue) edge. We say that the new edge $e$ is positive if $t(e)$ is connected in $G$ by a red edge to $L_{n-1}$ and $h(e)$ is connected to $L_{n}$. Otherwise, $t(e)$ is connected to $L_{n}$ and $h(e)$ to $L_{n-1}$ and we say that $e$ is negative.
Figure 2: The cycle $C$ with new edges depicted in blue, along with their sign, and old paths existing in $G$ depicted by a black dotted line. (a) All new edges have the same sign, $P$ is some old path connecting adjacent blue edges, (b) Not all new edges have the same sign, there are at least two old paths ($P_{1}$ and $P_{2}$ in this case) connecting different sign blue edges.
If all new edges in $C$ have the same sign, as seen in Figure 2 (a), there is a path $P$ in $C$ of length at most k1k-1 that exists in $G$ and connects two same sign blue edges. Therefore, $P$ with two red edges forms a path connecting $L_{n}$ with $L_{n-1}$. It follows that $d_{G}(v^{\prime},u^{\prime})\leq 2\mbox{diam}(T)+k-1+2\leq 2\mbox{diam}(T)+g=D_{0}$, contradicting our assumption that $\mbox{dist}_{G}(v^{\prime},u^{\prime})>D_{0}$.
Otherwise, not all new edges have the same sign, as seen in Figure 2 (b). Therefore, there are at least two paths $P_{1}$ and $P_{2}$ in $C$, present in the original graph $G$, connecting two different sign new edges. Choosing $P$ to be the shortest of the two paths, yields a path $P$ of length at most $\lfloor(k-2)/2\rfloor$, with endpoints $h(e)$ and $t(f)$ for new edges $e,f$. Furthermore, both $h(e)$ and $t(f)$ are connected by red edges in $G$ to the same layer, either $L_{n}$ or $L_{n-1}$. Therefore, $G$ has a short cycle consisting of $P$, two red edges and a traversal of the tree $T$. The length of this cycle is at most $\lfloor(k-2)/2\rfloor+2+\mbox{diam}(T)\leq k/2+1+\mbox{diam}(T)<g/2+1+\mbox{diam}(T)$, which is less equal than $g$ for $g\geq g_{0}$, yielding a contradiction. ∎
Lemma 3.4.
For any undirected connected graph $H$ with minimal degree at least two and any $g\geq 1$, there exists a covering graph $G$ with $\mbox{girth}(G)\geq g$.
@@
Assume first that $H$ does not have half-loops. The proof is by induction on $g$, where the base $g=1$ trivially holds for $G=H$. By induction, assume that $G$ is a graph covering $H$ with $\mbox{girth}(G)\geq g-1$, where the number of length g1g-1 cycles $\phi(G)$ is as small as possible. If $\phi(G)=0$, then $\mbox{girth}(G)\geq g$, and we are done. Otherwise, assume $\phi(G)>0$ and let $G_{2}$ be a random 2-lift of $G$. Namely, a random graph with vertex set $V(G)\times\{0,1\}$, where for each edge $e\in\overline{E}(G)$ we pick the permutation $\mbox{perm}_{\pi}(e)\in S_{2}$ independently at random such that $\mbox{perm}_{\pi}(e)=\mbox{perm}_{\pi}(e^{-1})$. We observe that all 2-lifts of $G$ have girth at least g1g-1, and argue that among them there is a 2-lift with a smaller number of length g1g-1 cycles. Indeed, as every length g1g-1 cycle $C$ in $G$ is turned into either one cycle of length $2(g-1)\geq g$ or two cycles of length g1g-1, with equal probabilities, depending on the product of $\mbox{perm}_{\pi}(e)$ for $e\in C$, it follows that the expected value of $\phi(G_{2})$ is $\phi(G)$. However, since the lift with $\mbox{perm}_{\pi}(e)=\text{id}$ for all $e$ has $2\phi(G)$ length g1g-1 cycles, it follows that there exists some 2-lift $G^{\prime}$ of $G$ with $\phi(G^{\prime})<\phi(G)$, yielding a contradiction.
If $H$ has half-loops, we let the graph $H^{\prime}$ be some 2-lift of $H$, where the lift of the half-loop $e$ with $h(e)=t(e)=v$ is a single edge connecting the two vertices covering $v$. As the graph $H^{\prime}$ covers $H$ and does not have any half-loops, the required result is obtained by applying the preceding proof to $H^{\prime}$. ∎
@@ 4 Three Generalizations for Non-regular Graphs
Respectively, $n(H,g)=n({\cal L}(H),g)$ is the smallest girth $g$ lift of $H$, $n({\cal U}(H),g)$ is the smallest girth $g$ graph with the same universal cover as $H$ and $n({\cal D}(H),g)$ is the smallest girth $g$ graph with the same degree distribution as $H$. Since ${\cal L}(H)\subseteq{\cal U}(H)\subseteq{\cal D}(H)$ it follows that:
@@
    $\frac{1}{|\overline{E}|}\mathbf{1}^{T}(B_{H})^{r}\mathbf{1}\geq\Lambda^{r}$ for all $r\geq 0$.
@@
This result is related to the $\rho$ based lower bound of Theorem 1.1 by the following proposition. In that work, the authors prove that $\rho(H)\geq\Lambda(H)$ and specify two equivalent conditions for equality to hold.
Proposition 4.3 (Eisner, Hoory, [4]).
For a finite undirected and connected graph $H$ with minimal degree at least two and maximal degree greater than two, we have $\rho(H)\geq\Lambda(H)$. Furthermore, the following three conditions are equivalent:
  1. 1.
    $\rho(H)=\Lambda(H)$
  2. 2.
    For every non-backtracking cycle $C$ the geometric average of vertex degrees minus one along the cycle is equal to $\Lambda$,
    $\left(\prod_{v\in C}(\deg(v)-1)\right)^{\frac{1}{|C|}}=\Lambda(H).$
  3. 3.
    For every non-backtracking path $P$, where all its internal vertices (if any) have degree two and its endpoints have degree larger than two,
    $\left(\prod_{v\in P}(\deg(v)-1)\right)^{\frac{1}{2|P|}}=\Lambda(H).$
It follows from Claim 4.2, Proposition 4.3, inequalities (7), (6) and Theorem 1.1 that:
@@
It should be observed that because all three graphs are non-regular, we have strict inequality $\Lambda(H)>\overline{d}(H)-1$, in agreement with Claim 4.2. Furthermore, $\rho(H)>\Lambda(H)$, in agreement with Proposition [4]:
@@ 5 Concrete Constructions and Bounds for the $H_{23}$ Graph
In this section we explore non-asymptotic upper and lower bounds on the value of $n(H,g)$, for the graph $H=H_{23}$ depicted in Figure 3 (b). We discuss this in the following subsections: The Moore lower bound, The Erdös-Sachs Upper Bound, Construction Algorithms, and Results.
@@ 5.1 The Moore Lower bound
The Moore lower bound for lifts is stated in Lemma 3.1, where the size of the ball centered around some tree vertex can be obtained by (4) and (5), with $g=2r+1$:
$n(H,g)\geq n_{0}(H,g)=\max_{v\in V(\tilde{H})}|B_{r}(v)|=1+\max_{v\in V(H)}\sum_{i=1}^{r}\delta_{v}^{T}\left(B_{H}\right)^{i-1}\mathbf{1},$ (10)
where $\delta_{v}$ is the indicator function of $t^{-1}(v)$. One can derive a similar formula for even girth.
5.2 The Erdös-Sachs Upper Bound
We compute the Erdös-Sachs upper bound for $n(H,g)$ using Corollary 3.3. As the spanning tree $T$ has diameter one, it follows that for all $g\geq g_{0}=4$:
$n(H,g)\leq\min_{v\in V(\tilde{H})}|B_{g+2}(v)|=1+\min_{v\in V(H)}\sum_{i=1}^{g+2}\delta_{v}^{T}\left(B_{H}\right)^{i-1}\mathbf{1}.$
@@ 5.4 Result Summary
6 Conclusion & Open Problems
In this work we suggest alternatives, applicable to non-regular graphs, for the well studied problem of finding $n(k,g)$, which is the smallest size of a $k$-regular girth $g$ graph. Given a fixed base graph $H$, we define: (i) $n({\cal L}(H),g)=n(H,g)$ as the smallest size of a girth $g$ graph covering $H$, (ii) $n({\cal U}(H),g)$ as the smallest size of a girth $g$ graph with the same universal cover as $H$, and (iii) $n({\cal D}(H),g)$ as the smallest size of a girth $g$ graph with the same degree distribution as $H$. It should be noted that the second and third definitions offer a true generalization of $n(k,g)$, by taking any $k$-regular graph as $H$.
We prove that the first two definitions are equivalent up to a multiplicative constant, and we show that the well known Moore lower bound and Erdös-Sachs upper bound hold in the new setting as well, with the exponent base being $\rho(H)$ rather than k1k-1. For the 3rd definition we have a Moore bound result by Alon, Hoory and Linial [1]. However, there is no Erdös-Sachs result and the Moore bound, with exponent base $\Lambda(H)$, might be far from achievable as we conjecture is the case for $H_{23}$.
@@ 6 Conclusion & Open Problems
    Conjecture: $n({\cal D}(H_{23}),g)\geq{\rho(H_{23})}^{(1+o(1))g/2}$
    Namely, the exponent base lower bound on the size of girth $g$ graphs where half the vertices have degree 2 and the other half have degree 3, can be improved from $\Lambda(H_{23})$ to $\rho(H_{23})$. An even bolder conjecture would be that graphs achieving the minimum cannot have adjacent degree 2 vertices.
@@
    One should note that by Proposition 4.3 states that $\rho(G)>\Lambda(H_{23})$ for any graph $G\in{\cal D}(H_{23})$. However, it falls short of proving that $\rho(G)>\Lambda(H_{23})+\epsilon$ for some explicit $\epsilon>0$.
@@
    Problem: Prove that any of the greedy algorithms listed in Section 5 satisfies the asymptotic Erdös-Sachs upper bound of Theorem 1.1 whp.
@@
    Problem: Prove that the Moore lower bound for $H_{23}$, given by (10) cannot hold as equality for large values of $g$, even when taking the divisibility constraint that $n(H_{23},g)$ must be divisible by four.
@@
    Problem: Provide a number theoretic construction of girth $g$ graphs covering some non-regular graph $H$ of size ${\rho(H)}^{(1+o(1))\cdot 3g/4}$.